4.4 For the enzyme reaction:


The Michaelis and Menten constant, Km is actually a summary of 3 terms:

(a) What are these? 

(b) What assumption does it make?

(c) Define and calculate  Km from the above equation given:

(d) Which is the rate-limiting step? And why?

MBBE 402 BiochemistryExamination L 2015 For the enzyme reaction: ki k2 The Michaelis and Menten constant, K is actually a summary o (a) What are these? Cata iy hc e fActency (b) What assumption does it make? (c) Define and calculate K.m from the above equation given: k.2- 11 SecM (d Which is the rate-limiting step? And why? (10 Marks) f 3 terms: S subshate Concen trah on Vo = Vmax [S] km [Si cat Ctumaver# Celr ↑kcat ↓h catalytic efftiercy b) If a reaction has serval steps and one is cleary Kcat l$12 to rate crnstrmi rate timlting
Answer #1

At steady state,

Rate of formation of ES = Rate of break down of ES

K1[E] [S] = (K-1+K2) [ES]

[E] [S] / [ES] = (K-1 + K2) / K1 = Km = Michaelis menten constant

On substituting,

Km = (2+11) / 8

Km = 1.625

Since it is a constant there is no unit.

Note: the unit given for K1 should be sec-1 m-1. The given unit is a typo.

Km shows the binding affinity of a substrate towards the enzyme. It gives the concentration of the substrate [S] when Vo = 1/2 Vm. (Vo = reaction velocity and Vm is maximum velocity)